# Spatial-Network Treatment Effects: A Continuous Functional Approach

### v3

@@ 2.8 Mutual Information Characterization
Theorem 2 (Mixed Effect as Mutual Information).
@@ 2.9 General Equilibrium Amplification
Proposition 8 (Amplification Factor).
@@ 3 Identification and Estimation
3.1 From Theory to Econometric Objects
The theoretical framework provides structural objects—the source term $S$, treatment functional $\tau$, and propagation parameters $(\nu_{s},\nu_{n},\kappa,\lambda)$—that must be connected to empirical counterparts.
The source term $S(\mathbf{x},t,\alpha)$ is directly observable from policy variation. For minimum wage policy, $S$ equals the dollar amount by which the new minimum wage exceeds the prevailing wage at each location-time-industry cell.
The treatment functional $\tau(\mathbf{x},t,\alpha)$ is not directly observed. We observe outcomes $Y$ that depend on $\tau$:
$Y_{i}=Y_{i}^{(0)}+\tau(\mathbf{x}_{i},t,\alpha_{i})+\varepsilon_{i}$ (51)
where $Y_{i}^{(0)}$ is the potential outcome without treatment. The identification challenge is recovering $\tau$ (or its parameters) from observable variation.
The connection to reduced-form coefficients proceeds through the regression specification:
$Y_{i}=\beta_{0}+\beta_{s}S_{i}+\beta_{n}\tilde{N}_{i}+\lambda S_{i}\cdot\tilde{N}_{i}+\mathbf{X}_{i}^{\prime}\boldsymbol{\gamma}+\varepsilon_{i}$ (52)
where $S_{i}$ is direct source exposure, $\tilde{N}_{i}=\sum_{j}G_{ij}S_{j}$ is network-weighted exposure to others’ sources, and the interaction captures the mixed effect.
3.2 The No-Spillover Benchmark
Before addressing the full model, we examine the benchmark case with no spillovers: $\nu_{s}=\nu_{n}=0$. This case corresponds to SUTVA holding and provides a foundation for understanding when spillovers matter.
3.2.1 The Model Without Spillovers
@@
This is an ordinary differential equation at each point, with no spatial or network coupling. The solution is:
@@
Treatment intensity at each location depends only on (i) the initial condition at that location, discounted by market adjustment, and (ii) the cumulative source at that location, discounted by market adjustment. There is no propagation from other locations.
@@
This is the continuous treatment effect model studied by Hirano and Imbens (2004) and Kennedy et al. (2017).
3.2.2 Identification Under No Spillovers
@@
Potential outcomes under any treatment level $s$ are independent of actual treatment assignment conditional on covariates. This is the continuous-treatment analog of selection on observables.
@@
The generalized propensity score (GPS), defined as the conditional density of treatment given covariates:
$r(s,\mathbf{x})=f_{S|X}(s\,|\,\mathbf{x})$ (57)
provides the basis for estimation. Under unconfoundedness, conditioning on the GPS removes selection bias:
$\mathbb{E}[Y\,|\,S=s,R=r]=\mu(s,r)$ (58)
where $\mu(s,r)$ is the dose-response surface.
The average dose-response function is then:
$\mu(s)=\mathbb{E}[\mu(s,r(s,X))]$ (59)
This approach, while valid under SUTVA, fails when spillovers are present. If treatment at location $j$ affects outcomes at location $i\neq j$, then $Y_{i}$ depends not just on $S_{i}$ but on the entire vector $(S_{1},\ldots,S_{N})$. The GPS conditions only on own treatment and own covariates, missing the spillover channel.
3.2.3 Testing for Spillovers
The no-spillover benchmark provides testable predictions. If $\nu_{s}=\nu_{n}=0$:
  1. 1.
    No spatial gradient: Treatment effects should not vary with distance from treated regions:
    $H_{0}:\frac{\partial\tau}{\partial d}=0$ (60)
    where $d$ is distance to nearest treated unit.
  2. 2.
    No network gradient: Treatment effects should not vary with network exposure:
    $H_{0}:\frac{\partial\tau}{\partial\tilde{N}}=0$ (61)
  3. 3.
    No interaction: The spatial gradient should not depend on network exposure:
    $H_{0}:\frac{\partial^{2}\tau}{\partial d\,\partial\tilde{N}}=0$ (62)
@@ 3.2.3 Testing for Spillovers
These tests can be implemented by augmenting the no-spillover specification with distance and network terms and testing their joint significance.
3.3 Identification Challenges with Spillovers
When spillovers are present ($\nu_{s}>0$ or $\nu_{n}>0$), identification confronts three fundamental challenges.
The Reflection Problem
The reflection problem, identified by Manski (1993), arises because observed correlations between an agent’s outcome and neighbors’ outcomes may reflect causal spillovers, correlated unobservables, or simultaneity. When California wages rise, Nevada wages may rise because of spillovers (causal), because both states face common shocks (confounding), or because Nevada wages also affect California (reverse causality).
Spatial Confounding
Spatial confounding arises because unobserved location-specific factors correlate with both treatment exposure and outcomes. Counties near California may have unobserved characteristics that jointly determine their wage levels and their proximity to California’s labor markets.
Network Endogeneity
Network endogeneity arises because network connections form endogenously. Firms connected through supply chains may be similar on unobserved dimensions that jointly determine their network position and their response to treatment.
3.4 Identification Strategy
We address these challenges through three complementary approaches.
3.4.1 Spatial Regression Discontinuity
The spatial RD design exploits sharp treatment boundaries at state borders, following Lee and Lemieux (2010) and Dube et al. (2010).
At state borders where minimum wage changes discontinuously, other factors affecting wages evolve smoothly. Let $d_{i}$ denote signed distance from unit $i$ to the nearest treatment boundary, with $d_{i}>0$ for treated units.
Assumption 1 (Continuity at Boundary).
Potential outcomes under control are continuous at the boundary:
$\lim_{d\to 0^{+}}\mathbb{E}[Y_{i}(0)|d_{i}=d]=\lim_{d\to 0^{-}}\mathbb{E}[Y_{i}(0)|d_{i}=d]$ (63)
Proposition 9 (Spatial RD Identification).
Under the continuity assumption:
The direct spatial effect is identified from the discontinuity:
$\beta_{s}=\lim_{d\to 0^{+}}\mathbb{E}[Y_{i}|d_{i}=d]-\lim_{d\to 0^{-}}\mathbb{E}[Y_{i}|d_{i}=d]$ (64)
The spatial decay parameter is identified from how the discontinuity attenuates:
$\kappa_{s}=-\frac{\partial}{\partial d}\log|\tau^{\text{RD}}(d)|$ (65)
3.4.2 Network Instrumental Variables
The network IV strategy uses predetermined network connections as instruments, following Bramoullé et al. (2009).
Historical network connections—supply chain relationships established before treatment—cannot respond to current shocks. Let $\mathbf{G}^{\text{hist}}$ denote historical connections and define the instrument:
$Z_{i}=\sum_{j\neq i}G_{ij}^{\text{hist}}S_{j}$ (66)
Assumption 2 (Exclusion Restriction).
Historical network connections affect current outcomes only through current network exposure:
$\mathbb{E}[Z_{i}\varepsilon_{i}]=0$ (67)
Assumption 3 (Relevance).
Historical and current networks are correlated:
$\text{Cov}(Z_{i},N_{i})\neq 0$ (68)
Proposition 10 (Network IV Identification).
Under the exclusion and relevance assumptions:
$\beta_{n}=\frac{\text{Cov}(Y_{i},Z_{i})}{\text{Cov}(N_{i},Z_{i})}$ (69)
3.4.3 Entropy-Based Moment Conditions
The entropy moments leverage theoretical predictions from Section 2 about treatment distribution dynamics.
The theoretical framework predicts exponential decay of relative entropy:
$D_{KL}(p_{t}\|p_{\infty})=D_{KL}(p_{0}\|p_{\infty})e^{-2\lambda_{2}t}$ (70)
where $\lambda_{2}$ is the spectral gap determined by $(\nu_{s},\nu_{n},\kappa)$.
This generates moment conditions:
$\mathbb{E}\left[\log\hat{D}_{KL}(t)+2\lambda_{2}t-c\right]=0$ (71)
Proposition 11 (Entropy Identification).
The spectral gap is identified from the time series of KL divergences:
$\lambda_{2}=-\frac{1}{2T}\log\frac{\hat{D}_{KL}(T)}{\hat{D}_{KL}(0)}$ (72)
The interaction coefficient is identified through its relationship to mutual information:
$\lambda=I(\mathbf{x};\alpha)$ (73)
@@ 3.5 Monte Carlo Evidence
We conduct comprehensive Monte Carlo simulations to assess the framework’s performance under various conditions.
3.5.1 Simulation Design
We simulate a spatial-network economy with $N=500$ units over $T=20$ periods. Units are located on a $25\times 20$ grid with coordinates $\mathbf{x}_{i}\in[0,100]^{2}$. Market positions $\alpha_{i}\sim\text{Uniform}[0,1]$ are drawn independently. The network adjacency matrix $G_{ij}$ is constructed with connection strength decaying exponentially with distance in $\alpha$: $G_{ij}=\exp(-|\alpha_{i}-\alpha_{j}|/0.2)$, normalized so rows sum to one.
We consider four simulation scenarios:
  1. 1.
    No spillovers: $\nu_{s}=\nu_{n}=0$, $\kappa=0.3$, $\lambda=0$
  2. 2.
    Spatial spillovers only: $\nu_{s}=2.0$, $\nu_{n}=0$, $\kappa=0.3$, $\lambda=0$
  3. 3.
    Network spillovers only: $\nu_{s}=0$, $\nu_{n}=0.5$, $\kappa=0.3$, $\lambda=0$
  4. 4.
    Full model: $\nu_{s}=2.0$, $\nu_{n}=0.5$, $\kappa=0.3$, $\lambda=0.4$
3.5.2 Estimators Compared
We compare five estimators:
  1. 1.
    TWFE: Two-way fixed effects regression
  2. 2.
    Continuous GPS: Generalized propensity score following Hirano and Imbens (2004)
  3. 3.
    Spatial lag: Spatial autoregressive model
  4. 4.
    No-spillover PDE: Our framework with $\nu_{s}=\nu_{n}=0$ imposed
  5. 5.
    Full PDE: Our framework estimating all parameters
3.5.3 Results: Binary Treatment Timing
@@
Notes: Event study plots from 1,000 Monte Carlo replications with binary treatment timing (treatment switches on at $t=5$ and remains constant). Panel A: No spillovers ($\nu_{s}=\nu_{n}=0$); all estimators perform similarly. Panel B: Full spillover model; TWFE and GPS underestimate by 25–35%, no-spillover PDE underestimates by 30%, only full PDE recovers true effect. Effects measured at border region for Panel B.

### v4

@@ 2.12 Mutual Information Characterization
Theorem 2.3 (Mixed Effect as Mutual Information).
@@ 2.13 General Equilibrium Amplification
Proposition 2.12 (Amplification Factor).
@@ 3 Identification and Estimation
3.1 The No-Spillover Benchmark and Conventional Treatment Effects
Before addressing the full spatial-network model, we examine the benchmark case with no spillovers: $\nu_{s}=\nu_{n}=0$. This case corresponds to standard treatment effect estimation where the Stable Unit Treatment Value Assumption (SUTVA) holds. Establishing this connection clarifies that our framework nests conventional methods rather than replacing them.
3.1.0.1 The Model Without Spillovers.
@@ 3.1 The No-Spillover Benchmark and Conventional Treatment Effects
This is an ordinary differential equation at each point $(\mathbf{x},\alpha)$, with no spatial or network coupling. The solution is:
@@
This is precisely the continuous treatment effect model studied by Hirano and Imbens (2004) and Kennedy et al. (2017).
3.1.0.2 Relationship to Generalized Propensity Score Methods.
@@
The generalized propensity score (GPS), defined as the conditional density of treatment given covariates $r(s,\mathbf{x})=f_{S|X}(s|\mathbf{x})$, provides the basis for estimation.
Proposition 3.1 (Equivalence to GPS Estimation).
When $\nu_{s}=\nu_{n}=0$, our estimator reduces to the generalized propensity score estimator of Hirano and Imbens (2004).
Proof.
Under (80), the dose-response function is:
$\mu(s)=\mathbb{E}[Y_{i}(s)]=\mathbb{E}[\mathbb{E}[Y_{i}|S_{i}=s,r(s,X_{i})]]$ (81)
This is identified by regression of $Y_{i}$ on $S_{i}$ and $r(S_{i},X_{i})$. When spatial and network terms are absent from our specification, we estimate exactly this regression. ∎
This proposition establishes that conventional GPS methods are a special case of our framework. When tests fail to reject the hypothesis that spillovers are absent, practitioners should use GPS methods, which are simpler to implement and communicate. The framework adds value only when spillovers are empirically present.
3.1.0.3 Relationship to Doubly Robust Estimation.
Recent work by Kennedy et al. (2017) develops doubly robust estimators for continuous treatment:
$\hat{\mu}(s)=\frac{1}{N}\sum_{i=1}^{N}\left[\frac{K_{h}(S_{i}-s)}{f_{S}(s)}Y_{i}-\frac{K_{h}(S_{i}-s)-f_{S|X}(s|X_{i})}{f_{S}(s)}\hat{\mu}(s,X_{i})\right]$ (82)
where $K_{h}$ is a kernel function and $\hat{\mu}(s,X_{i})$ is an outcome model. Our framework complements this by providing a structural model for $\mu(s,X_{i})$ through the master equation. When $\nu_{s}=\nu_{n}=0$:
$\mu(s,X_{i})=\frac{s}{\kappa}+X_{i}^{\prime}\gamma$ (83)
The doubly robust estimator is consistent if either the propensity score model OR the outcome model is correctly specified. Our approach provides a theoretically grounded outcome model from economic primitives, potentially improving finite-sample performance when the structural model is approximately correct.
3.1.0.4 Relationship to Difference-in-Differences.
For binary treatment timing (treatment switches on at time $t_{i}^{*}$), our framework connects to difference-in-differences (DiD):
$Y_{it}=\alpha_{i}+\gamma_{t}+\beta\cdot D_{it}+\varepsilon_{it}$ (84)
where $D_{it}=\mathbf{1}\{t\geq t_{i}^{*}\}$.
Setting $S_{it}=D_{it}$ and imposing $\nu_{s}=\nu_{n}=0$:
$\tau_{it}=\frac{1-e^{-\kappa(t-t_{i}^{*})}}{\kappa}\cdot S_{i}\cdot\mathbf{1}\{t\geq t_{i}^{*}\}$ (85)
For large $t-t_{i}^{*}$ (long-run effects), $\tau_{it}\to S_{i}/\kappa$, which is the standard DiD coefficient $\beta$. The continuous framework reveals that standard DiD estimates the long-run equilibrium effect, not the dynamic adjustment path. Recent heterogeneity-robust DiD estimators by Callaway and Sant’Anna (2021), Sun and Abraham (2021), and Borusyak et al. (2024) allow for dynamic effects. Our framework shows these dynamics should follow exponential decay at rate $\kappa$, providing a testable restriction beyond what atheoretical flexible specifications impose.
3.1.0.5 Testing for Spillovers.
The no-spillover benchmark provides three testable predictions. First, treatment effects should not vary with distance from treated regions, formally stated as testing whether $\partial\tau/\partial d=0$ where $d$ is distance to the nearest treated unit. Rejection indicates spatial spillovers with $\nu_{s}>0$. Second, treatment effects should not vary with network exposure, tested as $\partial\tau/\partial\tilde{N}=0$ where $\tilde{N}=\sum_{j}G_{ij}S_{j}$ is network-weighted treatment exposure. Rejection indicates network spillovers with $\nu_{n}>0$. Third, the spatial gradient should not depend on network exposure, tested as $\partial^{2}\tau/\partial d\,\partial\tilde{N}=0$. Rejection indicates mixed effects with $\lambda>0$.
@@
These tests can be implemented by augmenting specification (79) with distance and network exposure terms:
$Y_{i}=\beta_{0}+\beta_{s}S_{i}+\beta_{d}f(d_{i})+\beta_{n}\tilde{N}_{i}+\beta_{\lambda}f(d_{i})\cdot\tilde{N}_{i}+\mathbf{X}_{i}^{\prime}\boldsymbol{\gamma}+\varepsilon_{i}$ (86)
Testing $H_{0}:\beta_{d}=\beta_{n}=\beta_{\lambda}=0$ jointly assesses whether spillovers are present. Importantly, this test does not assume the PDE structure, relying only on correlations between outcomes and spatial-network exposure measures. The nesting structure creates a one-sided risk profile: when spillovers are absent, our framework correctly detects this and produces estimates equivalent to conventional methods. When spillovers are present, our framework captures them while conventional methods that maintain SUTVA fail, leading to biased estimates and invalid inference.
@@ 3.2 Monte Carlo Evidence
This subsection presents Monte Carlo evidence on the finite-sample properties of our estimator compared to conventional methods. The simulations demonstrate three key results. First, conventional estimators exhibit substantial bias when spillovers are present, with bias magnitudes of 25–38% depending on the method. Second, our framework maintains correct inference across all configurations, including when spillovers are absent. Third, this creates a one-sided risk profile: when researchers are uncertain about whether spillovers exist, our framework provides valid inference regardless, while conventional methods risk severe bias if spillovers are incorrectly ignored.
3.2.0.1 Simulation Design.
We simulate spatial-network treatment effect data matching the empirical setting of U.S. minimum wage policy. The simulation design follows the theoretical framework developed in Section 2 while introducing realistic features of empirical applications.
3.2.0.2 Geographic and Network Structure.
We construct a stylized geography with $N=500$ units distributed over a two-dimensional spatial domain $0,100]\times[0,100]$ measured in miles. Units are randomly located with a spatial Poisson process ensuring realistic clustering patterns observed in actual county distributions. Each unit $i$ has spatial coordinates $\mathbf{x}_{i}=(x_{i}^{(1)},x_{i}^{(2)})$ and a market position coordinate $\alpha_{i}\in[0,1]$ representing position in production networks.
The network structure captures supply chain connections. We generate network links using a gravity model where the probability that unit $i$ connects to unit $j$ is:
$p_{ij}=\exp\left(-\theta_{d}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|-\theta_{\alpha}|\alpha_{i}-\alpha_{j}|\right)$ (87)
The parameters $(\theta_{d},\theta_{\alpha})$ control how network connections decay with geographic and market distance. We set $\theta_{d}=0.02$ per mile and $\theta_{\alpha}=2$ to generate networks where firms preferentially connect to both nearby firms and firms at similar production stages. This creates realistic industry clustering patterns where geographic proximity and supply chain position are correlated, generating positive mutual information $I(\mathbf{x};\alpha)>0$ and thereby spatial-network interaction effects.
The average degree is $\mathbb{E}[\text{degree}_{i}]\approx 15$ connections per firm, matching empirical supply chain data from the Bureau of Economic Analysis input-output tables. The network exhibits realistic features including clustering (average clustering coefficient 0.42), small-world properties (average path length 3.2), and degree heterogeneity (coefficient of variation 0.78).
3.2.0.3 Data Generating Process.
Outcomes are generated from the continuous master equation with added measurement error:
$Y_{i}=\tau(\mathbf{x}_{i},\alpha_{i})+\mathbf{X}_{i}^{\prime}\boldsymbol{\gamma}+\varepsilon_{i}$ (88)
The treatment functional $\tau(\mathbf{x}_{i},\alpha_{i})$ satisfies:
$0=\nu_{s}\nabla^{2}\tau+\nu_{n}\frac{\partial^{2}\tau}{\partial\alpha^{2}}-\kappa\tau+\lambda\frac{\partial\tau}{\partial x^{(1)}}\frac{\partial\tau}{\partial\alpha}+S(\mathbf{x}_{i},\alpha_{i})$ (89)
We solve this PDE numerically using a finite element method on a fine grid and interpolate to obtain $\tau(\mathbf{x}_{i},\alpha_{i})$ at each simulated unit location. This ensures the data generating process exactly matches the theoretical framework.
The source term represents policy treatment:
$S_{i}=S_{0}\cdot\mathbf{1}\{x_{i}^{(1)}>50\}\cdot(1+0.3\cdot\alpha_{i})$ (90)
This mimics a minimum wage increase in the eastern half of the domain (states on the right side), where treatment intensity varies with industry position $\alpha_{i}$. Lower-wage industries ($\alpha_{i}$ near 0) receive larger proportional increases. We set $S_{0}=0.10$, representing a 10 cent minimum wage increase.
Control variables include:
$X_{i,1}$ $=0.5\alpha_{i}+u_{i,1}\qquad$ (industry characteristics) (91)
$X_{i,2}$ $=\log(\text{distance to border}_{i}+1)+u_{i,2}\qquad$ (geography) (92)
$X_{i,3}$ $=\text{degree}_{i}/15+u_{i,3}\qquad$ (network centrality) (93)
where $u_{ij}\sim N(0,0.25^{2})$ are idiosyncratic shocks. These controls introduce realistic confounding: outcomes correlate with industry characteristics, geographic position, and network centrality even absent treatment effects. The error term is $\varepsilon_{i}\sim N(0,\sigma_{\varepsilon}^{2})$ with $\sigma_{\varepsilon}=0.05$, yielding a signal-to-noise ratio matching typical empirical applications.
3.2.0.4 Parameter Configurations.
We simulate under four configurations varying the spillover parameters:
Configuration 1 (No Spillovers).
$\nu_{s}$ $=0,\qquad\nu_{n}=0,\qquad\lambda=0,\qquad\kappa=0.25$
This is the SUTVA case where conventional methods should be valid.
Configuration 2 (Spatial Spillovers Only).
$\nu_{s}$ $=100\text{ sq mi/Q},\qquad\nu_{n}=0,\qquad\lambda=0,\qquad\kappa=0.25$
Spillovers propagate geographically but not through networks.
Configuration 3 (Network Spillovers Only).
$\nu_{s}$ $=0,\qquad\nu_{n}=0.015,\qquad\lambda=0,\qquad\kappa=0.25$
Spillovers propagate through supply chains but not geographically.
Configuration 4 (Full Spatial-Network Model).
$\nu_{s}$ $=100\text{ sq mi/Q},\qquad\nu_{n}=0.015,\qquad\lambda=0.04\text{ nats},\qquad\kappa=0.25$
Both channels present with significant interaction. These parameters match empirical estimates from Section 4. For each configuration, we generate $M=1000$ Monte Carlo replications.
3.2.0.5 Estimators Compared.
We compare six estimators spanning conventional and spatial-network approaches:
Method 1 (Two-Way Fixed Effects (TWFE)).
The standard panel regression with unit and time fixed effects:
$Y_{it}=\alpha_{i}+\gamma_{t}+\beta\cdot S_{it}+\mathbf{X}_{it}^{\prime}\boldsymbol{\gamma}+\varepsilon_{it}$ (94)
We use our cross-sectional data by treating spatial location bins as ”time periods,” yielding a pseudo-panel structure. This represents the most common approach in applied work. Standard errors are clustered by geographic region.
Method 2 (Difference-in-Differences (DiD)).
We implement heterogeneity-robust DiD following Callaway and Sant’Anna (2021):
$\hat{\tau}_{\text{DiD}}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\,|\,D_{i}=1]-\mathbb{E}[Y_{i}(1)-Y_{i}(0)\,|\,D_{i}=0]$ (95)
Treatment group is $D_{i}=\mathbf{1}\{x_{i}^{(1)}>50\}$, control group is $D_{i}=0$. We use inverse probability weighting with GPS to account for confounders. This represents best-practice DiD accounting for treatment effect heterogeneity.
Method 3 (Generalized Propensity Score (GPS)).
Following Hirano and Imbens (2004), we estimate:
$\hat{\tau}_{\text{GPS}}(s)=\mathbb{E}\left[\frac{K_{h}(S_{i}-s)}{f_{S}(s)}Y_{i}\,\bigg|\,r(s,X_{i})\right]$ (96)
where $r(s,X_{i})$ is the generalized propensity score. We use kernel smoothing with bandwidth selected by cross-validation. This is the leading method for continuous treatment effects under SUTVA.
Method 4 (Spatial Regression Discontinuity (Spatial RD)).
We exploit the sharp treatment boundary at $x^{(1)}=50$:
$\hat{\tau}_{\text{RD}}=\lim_{x^{(1)}\downarrow 50}\mathbb{E}[Y_{i}|x_{i}^{(1)}]-\lim_{x^{(1)}\uparrow 50}\mathbb{E}[Y_{i}|x_{i}^{(1)}]$ (97)
We use local linear regression with triangular kernel and bandwidth selected by Imbens and Kalyanaraman (2012) optimal bandwidth algorithm. This accounts for spatial spillovers at the border but not network spillovers or interaction effects.
Method 5 (Network IV).
We use lagged network structure as instruments for current network exposure:
$\text{First stage:}\qquad\tilde{N}_{i}$ $=\sum_{j}G_{ij}^{(t-1)}S_{j}+\mathbf{X}_{i}^{\prime}\boldsymbol{\pi}+\eta_{i}$ (98)
$\text{Second stage:}\qquad Y_{i}$ $=\beta_{0}+\beta_{s}S_{i}+\beta_{n}\hat{\tilde{N}}_{i}+\mathbf{X}_{i}^{\prime}\boldsymbol{\gamma}+\varepsilon_{i}$ (99)
where $G_{ij}^{(t-1)}$ is predetermined network structure. This accounts for network spillovers but not spatial spillovers or interaction effects.
Method 6 (Full Spatial-Network GMM (Our Framework)).
We implement the complete GMM estimator combining spatial RD, network IV, and entropy moment conditions:
$\mathbb{E}[m_{\text{RD}}(\boldsymbol{\theta})]$ $=0\qquad\text{(spatial discontinuity)}$ (100)
$\mathbb{E}[m_{\text{IV}}(\boldsymbol{\theta})]$ $=0\qquad\text{(network instruments)}$ (101)
$\mathbb{E}[m_{\text{entropy}}(\boldsymbol{\theta})]$ $=0\qquad\text{(distributional dynamics)}$ (102)
Standard errors use spatial-network HAC accounting for both dependence sources. The estimator identifies $(\nu_{s},\nu_{n},\lambda,\kappa)$ jointly.
3.2.0.6 Results: Bias and Efficiency.
Table 3 reports bias and root mean squared error (RMSE) for each estimator across the four configurations. The results provide stark evidence for the one-sided risk profile property.
Table 3: Monte Carlo Results: Bias and RMSE
Config 1: No Config 2: Spatial Config 3: Network Config 4: Full
Spillovers Only Only Model
Estimator Bias RMSE Bias RMSE Bias RMSE Bias RMSE
Panel A: Direct Effect Estimate ($\beta_{s}$, true value = 0.100)
TWFE 0.002 0.018 -0.031 0.042 -0.025 0.036 -0.038 0.051
DiD 0.001 0.016 -0.028 0.039 -0.023 0.034 -0.035 0.048
GPS -0.001 0.015 -0.029 0.040 -0.024 0.035 -0.036 0.049
Spatial RD 0.003 0.019 0.004 0.021 -0.024 0.035 0.002 0.020
Network IV 0.002 0.017 -0.030 0.041 0.003 0.019 -0.012 0.025
Full GMM 0.001 0.016 0.002 0.018 0.001 0.017 0.003 0.019
Panel B: Total Effect at Border ($\tau_{\text{border}}$, varies by configuration)
(true = 0.100) (true = 0.153) (true = 0.148) (true = 0.171)
TWFE 0.002 0.018 -0.063 0.074 -0.058 0.069 -0.081 0.094
DiD 0.001 0.016 -0.059 0.071 -0.054 0.066 -0.076 0.089
GPS -0.001 0.015 -0.061 0.072 -0.056 0.067 -0.078 0.091
Spatial RD 0.003 0.019 0.006 0.022 -0.055 0.066 -0.021 0.034
Network IV 0.002 0.017 -0.062 0.073 0.005 0.021 -0.037 0.048
Full GMM 0.001 0.016 0.004 0.019 0.003 0.018 0.005 0.021
In Configuration 1 (no spillovers), all estimators are approximately unbiased with similar RMSE around 0.015–0.019. This confirms that when SUTVA holds, conventional methods work well and our framework adds no bias. The Full GMM estimator has RMSE of 0.016, essentially identical to GPS (RMSE = 0.015), showing our framework does not sacrifice efficiency when spillovers are absent.
In Configuration 2 (spatial spillovers only), conventional estimators ignoring spillovers—TWFE, DiD, GPS—exhibit bias of -0.028 to -0.031 for the direct effect, representing 28–31% bias relative to the true value of 0.100. This occurs because these methods attribute spillover effects to direct effects, biasing direct effect estimates downward. For total effects at the border (true value 0.153), the bias is even larger at -0.059 to -0.063, representing 38–41% bias. The Spatial RD estimator correctly recovers both direct and total effects (bias 0.004 and 0.006), while Network IV remains biased. The Full GMM estimator is essentially unbiased (bias 0.002 and 0.004).
Configuration 3 (network spillovers only) exhibits symmetric patterns. Conventional methods show bias of -0.023 to -0.025 for direct effects and -0.054 to -0.058 for total effects. Network IV correctly recovers effects while Spatial RD remains biased. Full GMM is again essentially unbiased.
Configuration 4 (full model with interaction) represents the most realistic and challenging case. Conventional methods exhibit the largest bias: -0.035 to -0.038 for direct effects (35–38%) and -0.076 to -0.081 for total effects (44–47%). Neither Spatial RD nor Network IV alone is sufficient, as each accounts for only one spillover channel. The Full GMM estimator maintains near-zero bias (0.003 and 0.005) by jointly estimating all parameters.
Table 4: Monte Carlo Results: Coverage Rates
Config 1: Config 2: Config 3: Config 4:
No Spillovers Spatial Only Network Only Full Model
Panel A: 95% Confidence Interval Coverage for Direct Effect
TWFE 0.946 0.587 0.652 0.524
DiD 0.951 0.612 0.678 0.548
GPS 0.948 0.598 0.665 0.537
Spatial RD 0.943 0.938 0.641 0.892
Network IV 0.947 0.593 0.945 0.814
Full GMM 0.949 0.947 0.951 0.946
Panel B: 95% Confidence Interval Coverage for Total Effect
TWFE 0.946 0.324 0.389 0.187
DiD 0.951 0.348 0.412 0.206
GPS 0.948 0.337 0.401 0.198
Spatial RD 0.943 0.921 0.378 0.698
Network IV 0.947 0.329 0.928 0.542
Full GMM 0.949 0.943 0.947 0.944
Table 4 examines confidence interval coverage rates. In Configuration 1 (no spillovers), all methods achieve nominal 95% coverage, confirming correct inference when SUTVA holds. However, when spillovers are present (Configurations 2–4), conventional methods suffer severe undercoverage. For direct effects in Configuration 4, TWFE achieves only 52.4% coverage—meaning confidence intervals miss the true parameter 48% of the time. For total effects, coverage drops to 18.7%, making conventional inference essentially uninformative.
The Full GMM estimator maintains 94.3–95.1% coverage across all configurations, demonstrating the one-sided risk profile: whether spillovers are absent or present, inference is valid. This robustness is crucial for applied work where researchers do not know a priori whether spillovers exist.
3.2.0.7 Results: Visual Evidence.
Figure 2 presents event study plots averaging across 1,000 Monte Carlo replications for each simulation scenario.
@@
Panel A shows results when the true DGP has no spillovers. All estimators perform similarly, tracking the true effect closely. The pre-treatment coefficients are centered at zero, confirming that all methods satisfy parallel trends when SUTVA holds. The no-spillover PDE and continuous GPS slightly outperform TWFE due to their use of continuous treatment intensity, but differences are modest.